Find x, m angle 1 & m angle 2, if ray BD bisects angle ABC and m angle 1=2x+8 and m angle ABC=50
Answer by mananth(16949)

BD bisects angle ABC
angle ABC = 50
1/2 angle = 25 deg
2x+8=25
2x=25-8
2x=17
x= 8.5
...
angle 1 = 2x+8 = 25
angle 2= 25
..
m.ananth@hotmail.ca